OpenStudy (anonymous):

help. how does i^18 = -1?

OpenStudy (alekos):

i^18 = (i^2)^9 = (-1)^9 = -1

OpenStudy (anonymous):

thank you! :)

OpenStudy (luigi0210):

Also: \[\LARGE i^1=\sqrt{-1}\] \[\LARGE i^2=-1\] \[\LARGE i^4=\sqrt{1}\] \[\LARGE i^3=-i\]

OpenStudy (luigi0210):

And as we see, 18 is a factor of 2, so it would be the i^2

OpenStudy (anonymous):

thanks for the help